OpenStudy (anonymous):

Select all the solutions of the given equation for 0° < θ < 360° secθsinθ - 2sinθ = 0 ans:60,90,120,180,240,300

myininaya (myininaya):

sin(x)(sec(x)-2)=0 Set both factors =0

myininaya (myininaya):

sin(x)=0 sec(x)=2 x=90 cos(x)=1/2 x=60, -60+360=300

OpenStudy (anonymous):

so its 90,60 and 300

OpenStudy (phi):

How surprising, but myi is confusing sin(x) with cos(x). sin is zero at 0 deg (excluded by your constraints 0<x< 360) and x=180 (and x=360, also excluded)

OpenStudy (mertsj):

So here is the bottom line. If sin x = 0 and x is between 0 and 360 degrees, then the angle x is 0 or 180 degrees. If the cos(x) = 1/2 and x is between ) and 360 degrees, then the angle is 60 and 300 degrees

OpenStudy (mertsj):

So choose 60, 180, and 300

myininaya (myininaya):

oops phi is right i made a mistake lol

myininaya (myininaya):

sin(x)=1 => x=90 my bad :(